If a nitric acid solution has a pH of 5, how does this acid concentration compare to the H+ contribution
of pure H2O, i.o.w. how many times bigger or smaller is it? Show calculations, concluding with a
proper sentence statement.
1
Expert's answer
2020-07-14T15:17:54-0400

pH = -log[H+]

Where [H+] is the concentration of H+ in Molarity,M

In Nitric acid the concentration of H+ is

[H+]= 10-5

The pH of pure water is 7

so in water the concentration of H+ ion is

[H+]= 10-7


So we can say that in Nitric acid the concentration 0f H+ ion is higher

and it is (10-5/10-7)= 100 times higher.